OpenStudy (anonymous):

Convert to rectangular r=6cot theta csc theta

OpenStudy (anonymous):

\[r=6 \cot \theta \csc \theta =6 \frac{ \cos \theta }{ \sin \theta }*\frac{ 1 }{ \sin \theta }=6\frac{ \cos \theta }{ \sin ^2\theta }\] \[r^2=6\frac{ r \cos \theta }{ \sin ^2 \theta }\] \[r^2 \sin ^2 \theta =6*r \cos \theta \] put\[ x =r \cos \theta ,y=r \sin \theta \] \[y^2=6x\]

OpenStudy (anonymous):

Isn't r^2=x^2+y^2

OpenStudy (anonymous):

What about (x-12)^2+(y+3)^2=144

OpenStudy (anonymous):

correct ,but\[r^2\sin ^2\theta =y^2\]

OpenStudy (anonymous):

Can you help with the other one

OpenStudy (anonymous):

what you want to do?

OpenStudy (anonymous):

it is a circle with centre (12,-3) and radius =12

OpenStudy (anonymous):

Convert to polar. I know it's a circle but how do I begin to turn it into polar if it doesn't have a center at the origin.

OpenStudy (anonymous):

just put \[x=r \cos \theta ,y=r \sin \theta \] and simplify